A rock is thrown straight down with initial velocity of 14, 5 m/s from a cliff . what is the rocks displacement after 2,0 s


1
Expert's answer
2022-04-17T17:06:27-0400

"d=v_0t+at^2/2\\ =14.5*2.0+9.8*2.0^2/2=48.6\:\rm m"